OpenStudy (anonymous):

find the quotient, to the nearest hundred, of 33,540 divided by 48.

OpenStudy (tkhunny):

\(\dfrac{33540}{48} = \dfrac{16770}{24} = \dfrac{8385}{12} = \dfrac{2795}{4} = \dfrac{69875}{100}\) I'm just kind of goofing off with this, trying to think of an interesting way to produce the result. Settign teh denominator to 100 seemed interesting. Of course, any calculator should slice through this with no problem.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

 can you finish it